I really need help on this
Find the equation of the line in slope intercept form given the following:
19.) m = 5 and point (3, -10)
20.) The line runs through (2, 3) and (-3, 7)

Answer by checkley71(8403)

Y=mX+b
-10=5*3+b
-10=15+b
b=-1)-15
b=-25 THEN THE EQUATION IS
Y=5X-25
-----------------------------------------------------
TO FIND THE SLOPE(m) WE MAKE A FRACTION (Y2-Y1/(X20-X1)
SLOPE=(7-3)/(-3-2)
SLOPE=4/-5
SLOPE=-4/5 NOW WE FIND THE Y INTERCEPT(b) BY USING ONE SET OF POINTS FOR X & Y
3=-4/5*2+b
b=3+4/5
b=19/5 THUS WE HAVE
Y=-4/5X+19/5 FOR THE LINE PASSING THROUGH THESE TWO POINTS.